[Ncep.nhc.nco_contacts] TAFB Shift Log -- Technical Edition -- Thu Jul 13, 2017

TAFB Shift Log nhc.tsbadmin at noaa.gov
Thu Jul 13 17:53:53 UTC 2017


TAFB Shift Log for Thu Jul 13 2017 13:53:49 GMT-0400 (EDT)

Shift: Day Shift

Pacific Forecaster: J Lewitsky

Atlantic Forecaster: M Nelson

Surface Analyst: J Aguirre-Echevarria

Technical Issues Summary:

GFE/NCP issues: None of the TAFB forecasters could log into the  
collaboration room or marine room at the start of the shift. Once we were  
able to, we were all kicked off. Josh @ NCF called to let us know that the  
NCF switched from BNCF to ANCF explaining the kick off. He also said that  
the collaboration server has many more problems when on BNCF (yesterday and  
the day before) than when on ANCF. He said that NHCN is particularly  
sensitive to the collaboration server as all it takes is for 1 packet out  
of 150 to drop out to result in problems. Josh said that he would look into  
the logs from this morning more and follow up with Chris Mello.

N-AWIPS issues:

Other issues: Ed informed us this morning that there were issues with the  
AC in perimeter areas. Temperature in TAFB ops area did rise for a while,  
but quickly cooled off later on.

Miscellaneous issues: TAFB PCs were not rebooted during the dayshift  
despite it being a Thursday.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170713/7222ff03/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list